# Multifidelity linear regression for scientific machine learning from scarce data

### v1

@@ 5 Conclusions and future work
We have introduced a multifidelity training strategy for linear regression problems which leverages data of varying fidelity and cost to improve the robustness of training to scarce high-fidelity data due to limited training budgets. Our approach is based on new multifidelity Monte Carlo estimators for the unknown linear regression coefficients, which we showed are unbiased with respect to the high-fidelity data distribution. We discussed multiple strategies for choosing the amount of data for each fidelity level, and for choosing the control variate coefficients in the multifidelity estimators, which can be viewed as hyperparameters of the method. We provide theoretical analyses showing that there are optimal choices of control variate coefficients when model statistics are known, and provide practical recommendations for choosing the hyperparameters when model statistics are unknown. Numerical experiments illustrate and verify our theoretical results and also demonstrate that the multifidelity training approach learns more robust linear regression models than a standard high-fidelity training approach when the training budget is limited and only scarce high-fidelity data are available.
@@
Many directions exist for future work. To make this multifidelity training approach more generally applicable in scientific machine learning, new approaches must be developed for linear regression models with vector outputs as well as for nonlinear regression models. These new methods will be critical for learning surrogate models that describe dynamics of time-dependent engineering systems rather than the input-to-scalar-output map we have focused on in this work. Even within the linear regression framework that we have introduced here, several avenues for further theoretical and practical investigation remain, including analysis of the method to determine optimal budget allocations tailored to the linear regression problem (instead of using the heuristic based on mean estimation that we recommend here), and exploring hybrid or active data acquisition strategies where some data are given but more can be obtained within budget constraints.
Acknowledgments. The authors are grateful to Florian Schäfer and Rob Webber for helpful discussions. EQ acknowledges support from the US Department of Energy, Office of Advanced Scientific Computing Research (ASCR), Applied Mathematics Program, award DE-SC0024721. AC acknowledges support from the US Department of Energy (DOE) grant number DE-SC0021239.

### v2

@@ 5 Conclusions and future work
We have introduced a multifidelity training strategy for linear regression problems which leverages data of varying fidelity and cost to improve the robustness of training to scarce high-fidelity data due to limited training budgets. Our approach is based on using multifidelity control variate estimators to reduce the variance of the linear regression estimators. We propose new multifidelity Monte Carlo estimators for the linear regression problem, which we show are unbiased with respect to the high-fidelity data distribution. We discussed multiple strategies for choosing the amount of data for each fidelity level, and for choosing the control variate coefficients in the multifidelity estimators, which can be viewed as hyperparameters of the method. We provide theoretical analyses showing that there are optimal choices of control variate coefficients when model statistics are known, and provide practical recommendations for choosing the hyperparameters when model statistics are unknown and must be estimated form limited samples. Numerical experiments illustrate our theoretical results and also demonstrate that the multifidelity training approach learns more accurate and robust linear regression models than the standard high-fidelity only training approach when the training budget is limited and only scarce high-fidelity data are available.
@@
Many directions exist for future work. Within the linear regression setting we have introduced here, several avenues for further theoretical and practical investigation remain, including analysis of the method to determine optimal model selection and budget allocations tailored to the linear regression problem (instead of using the heuristic based on mean estimation that we recommend here), and exploring hybrid or active data acquisition strategies where some data are given but more can be obtained within budget constraints. Detailed analysis and experiments comparing the MFMC-based estimator we have presented and alternatives based on ACV and MLBLUE would also be worthwhile. There are also many interesting questions to explore in using multifidelity control variates beyond the linear regression setting to more complex models, including neural networks. While some existing works have considered aspects of these problems [29, 15, 24], many theoretical and methodological questions remain.
Acknowledgments. The authors are grateful to Florian Schäfer and Rob Webber for helpful discussions, and to the anonymous referees for helpful feedback that has improved the clarity and completeness of the work. Work by EQ was supported by the US Department of Energy Office of Science Energy Earthshot Initiative as part of the ‘Learning reduced models under extreme data conditions for design and rapid decision-making in complex systems’ project under award number DE-SC0024721. AC acknowledges support from the US Department of Energy (DOE) grant number DE-SC0021239. AC and VS acknowledge support from DARPA Automating Scientific Knowledge Extraction and Modeling (ASKEM) program under Contract No. HR0011262087 (award number 653002). The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.
Appendix A Alternative control variate estimators for the linear regression problem
The main idea of this work is to propose a multifidelity approach to training linear regression models that combines high- and low-fidelity data using control variate estimators, thereby reducing the variance of the learned model predictions and increasing accuracy and robustness. In the main text, our focus is on the multifidelity learned model (12), which we describe as covariance estimation since the definitions of $\hat{\beta}^{\rm MF}$ and the associated learned model follow from estimates of $\hat{c}_{XY}^{\rm MF}$, as summarized here:
$\hat{f}^{\rm MF}(z)=\hat{f}(z;\hat{\beta}^{\rm MF}),\quad\hat{\beta}^{\rm MF}=C_{XX}^{-1}\hat{c}_{XY}^{\rm MF},$
and $\hat{c}_{XY}^{\rm MF}$ is given by (11). One could, however, directly formulate a control variate estimator for the unknown parameters $\beta$ and use this to define a model. We describe such an approach as parameter estimation:
$\hat{f}^{\rm MF}_{\rm param}(z)$ $=\hat{f}(z;\hat{\beta}^{\rm MF}_{\rm param}),\qquad\hat{\beta}^{\rm MF}_{\rm param}=\hat{\beta}_{m_{1}}^{(1)}+\tilde{A}(\hat{\beta}_{m_{2}}^{(2)}-\hat{\beta}_{m_{1}}^{(2)}),$ (23)
where $\hat{\beta}_{m_{i}}^{(j)}=\frac{1}{m_{i}}C_{XX}^{-1}X_{m_{i}}Y_{m_{i}}^{(j)}$ denotes the parameter estimator obtained from $m_{i}$ samples of data from model $j$, and $\tilde{A}\in\mathbb{R}^{d\times d}$. Finally, one could directly formulate a control variate estimator for the unknown model itself: we describe this approach as model estimation:
$\hat{f}^{\rm MF}_{\rm model}=\hat{f}(z;\hat{\beta}_{m_{1}}^{(1)})+\tilde{\alpha}\left(\hat{f}(z;\hat{\beta}_{m_{2}}^{(2)})-\hat{f}(z;\hat{\beta}_{m_{1}}^{(2)})\right),$ (24)
where we note $\tilde{\alpha}\in\mathbb{R}$ is a scalar since the output of $\hat{f}$ is one-dimensional.
The covariance, parameter, and model estimation approaches are closely related and are equivalent in some cases. In particular, if we let $A=\tilde{A}=\tilde{\alpha}I$, then all three estimators are the same. Additionally, the parameter estimation approach (23) is equivalent to the covariance approach if $\tilde{A}=C_{XX}^{-1}AC_{XX}$; that is, the parameter estimation approach represents a change of coordinates (via $C_{XX}^{-1}$) of the covariance estimation approach which is the focus of the main text.
We note that the model (24) and parameter (23) estimation approaches can also be analyzed in a similar vein to our analysis of Section 3.3. The optimal matrix coefficient for the parameter estimator (23) will be related to the optimal matrix coefficient for the covariance estimator by a similarity transform as discussed above, and will thus yield the same variance of parameters and model predictions. However, the optimal scalar coefficients for the three estimation approaches will generally be different, and will depend on second-order statistics of $C_{XX}^{-1}x(Z)f^{(i)}(Z)$ in the parameter case and of $x^{\top}(Z)C_{XX}^{-1}x(Z)f^{(i)}(Z)$ in the model case (recall the optimal coefficients depend on second-order statistics of $x(Z)f^{(i)}(Z)$ for the covariance case). When $C_{XX}$ is known exactly, the optimal coefficient for parameter estimation is as computable as the optimal coefficient for covariance estimation. However, when one generalizes to needing to estimate $C_{XX}$ as well, estimating the required second-order statistics for parameter estimation becomes more complicated. Estimating second-order statistics for the model estimation case is also generally more difficult than in the covariance estimation case: the model estimation setting essentially requires second-order statistics for the exact regression models themselves. We therefore have restricted our focus to the covariance estimation case, noting that further numerical and theoretical exploration of the alternative estimators discussed in this appendix is an area for future work.